a pie crust recipe calls for two cups of butter for each three cups of flour. Elliete is making enough crusts to use up eight cups of flour. How many cups of butter will she need?
Answer by Theo(13342)

the ratio of cups of butter to cups of flour = 2/3.
let x = the number of cups of butter to cups of flour.
when she used 8 cups of flour, you get:
2/3 = x/8
cross multiply to get 3x = 16
solve for x to get:
x = 16/3 cups of butter.
that's 5 and 1/3 cups of butter.

confirm that the ratio of cups of butter to cups of flour has been preserved by taking (16/3) and dividing it by 8 to 16 / 24 = 2/3.